OpenStudy (anonymous):

What is 8.75% written as a fraction in lowest terms?

OpenStudy (anonymous):

Answer: 7/80 Make whole numbers by multiplying numerator and denominator by 100: 8.75% = 8.75/100 = 875/10000 Then reduce: 875/10000 reduced by common divisor 1/125 = 7/80ths.